Commander Report – December 6th

Sol: 9
Summary Title: EVAs #12 and #13
Author’s name: Dr. Kristine Ferrone, Commander
Mission Status: Nominal
Commander’s Report:

Crew 269 experienced another day of great weather on Mars with mostly mild temperatures, sunny skies, and light high-altitude clouds, with morning to midday cloud cover and a sprinkle of rain. The crew awoke early this morning and enjoyed a quick breakfast of cereal, oatmeal, and dried fruit. After a morning planning meeting, XO, HSO, and Technology Officer conducted the morning EVA (#12) for the purpose of flying the drone and to continue demonstrating the VR/AR headset in the area around the Hab and along Cow Dung Rd. to Marble Ritual. During the morning EVA, Engineer and Green Hab Officer exchanged the furnace filter and changed the battery on space suit #6. Upon return of the EVA #12 crew, and after a quick lunch, CDR, Engineer, and Green Hab Officer embarked on EVA #13 south on Cow Dung Rd. and then on foot on Rte 1101 to Barrainca Butte for the purpose of regolith sample collection, radiation environment measurements, and handheld ham radio comm checks. Throughout the day, the crew was updated on the status of the power system by Tier 1 support and entered various levels of power conservation as directed. In the afternoon, HSO collected health stats, crew completed exercise protocols, and conducted a test of the weather balloon payload scheduled to launch later this week. For dinner the crew enjoyed pesto pasta with vegetables with homemade sourdough bread.

Copyright © The Mars Society. All rights reserved. | Main Site